OpenStudy (anonymous):

Solve the system of equations using elimination. 10m + 16n = 140 5m – 8n = 60

OpenStudy (luigi0210):

multiply the bottom by 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont get it

OpenStudy (luigi0210):

10m+16n=140 10m-16n=120 ------------------------------------------ 20m=260 m=13 ----------------------------------------- 10(13)+16n=140 130+16n=140 16n=10 n=5/8

OpenStudy (anonymous):

o u subtracted the bottom